# Framing global structural identifiability in terms of parameter symmetries

### v1

@@ 3.2.1 Analysing the structural identifiability of a toy model using parameter symmetries
This toy example illustrates that the standard differential algebra approach and the approach based on parameter symmetries arrive at the same conclusions regarding the identifiable parameter quantities. Additionally, the symmetry-based approach also yields the family of parameter symmetries of the toy model corresponding to the parameter transformations that preserve the observed outputs.
Subsequently, we show that both of these conclusions are also true in the general case, and we begin by providing an explanation for why the standard differential algebra approach always finds universal parameter invariants.
We demonstrate that the the standard differential algebra approach for analysing structural identifiability (Algorithm 1) will always find universal parameter invariants (Thm. 1 and Cor. 1). To this end, consider a system of output ODEs in Eq. (2) where the function $\Delta$ is a vector-valued multivariate polynomial for which the monomials are composed of the outputs $\mathbf{y}$ and derivatives of the outputs. In this case, the standard differential algebra approach is conducted in two steps where we first collect all coefficients of the monomials, and then we simplify these coefficients algebraically as well as reducing the number of coefficients as much as possible. This procedure yields all identifiable parameter quantities, and here we show that these identifiable parameter quantities are always given by universal parameter invariants in accordance with our definition of global structural identifiability (Cor. 1). This fact is a consequence of two fundamental properties of differential invariants of Lie symmetries.
Let $X_{\boldsymbol{\theta}}$ generate a parameter symmetry with $p$ parameters and denote its parameter invariants by $I_{1},\ldots,I_{\tilde{p}}$ where the number of invariants is an integer $\tilde{p}\in\{2,\ldots,p\}$. Then any differentiable function $F$ of these invariants denoted by $F\left(I_{1},\ldots,I_{\tilde{p}}\right)$ is itself a differential invariant.
Proof.
Thus, we have
$X_{\boldsymbol{\theta}}\left(F\left(I_{1},\ldots,I_{\tilde{p}}\right)\right)$ $=\sum_{\ell=1}^{p}\chi_{\ell}\frac{\partial F}{\partial\theta_{\ell}}=\sum_{\ell=1}^{p}\chi_{\ell}\sum_{j=1}^{\tilde{p}}\frac{\partial I_{j}}{\partial\theta_{\ell}}\frac{\partial F}{\partial I_{j}}=\sum_{\ell=1}^{p}\sum_{j=1}^{\tilde{p}}\chi_{\ell}\frac{\partial I_{j}}{\partial\theta_{\ell}}\frac{\partial F}{\partial I_{j}}$
$=\sum_{j=1}^{\tilde{p}}\frac{\partial F}{\partial I_{j}}\underset{=X_{\boldsymbol{\theta}}(I_{j})}{\underbrace{\left(\sum_{\ell=1}^{p}\chi_{\ell}\frac{\partial I_{j}}{\partial\theta_{\ell}}\right)}}=\sum_{j=1}^{\tilde{p}}\frac{\partial F}{\partial I_{j}}\underset{=0}{\underbrace{X_{\boldsymbol{\theta}}(I_{j})}}=0,$ (48)
which is the desired result. ∎
As a short detour, we clarify the difference between parameter invariants (Defn. 22) and universal parameter invariants (Defn. 5) using this fundamental property of differential invariants. Previously, we saw that the trivial parameter symmetry $\Gamma^{\boldsymbol{\theta}}_{0}:\boldsymbol{\theta}\mapsto\boldsymbol{\theta}$ is always a parameter symmetry of all possible models, and importantly all parameters $\boldsymbol{\theta}$ are parameter invariants of $\Gamma_{0}^{\boldsymbol{\theta}}$. Now, take another parameter symmetry $\Gamma_{\varepsilon}^{\boldsymbol{\theta}}$ which has a parameter invariant $I=F(\boldsymbol{\theta})$ defined by some non-constant, non-linear, multivariate and arbitrary function $F$. Moreover, let us assume that $I=F(\boldsymbol{\theta})$ is a universal parameter invariant. Then, the parameter invariant $I=F(\boldsymbol{\theta})$ of $\Gamma_{\varepsilon}^{\boldsymbol{\theta}}$ is also a parameter invariant of the trivial parameter symmetry $\Gamma_{0}^{\boldsymbol{\theta}}$ since $I$ is a function of the parameter invariants $\boldsymbol{\theta}$ of $\Gamma_{0}^{\boldsymbol{\theta}}$ (Prop. 1). For the same reason, all parameter invariants of $\Gamma_{\varepsilon}^{\boldsymbol{\theta}}$ are necessarily parameter invariants of $\Gamma_{0}^{\boldsymbol{\theta}}$. Nevertheless, the converse statement is not true as many of the parameter invariants of the trivial symmetry $\Gamma_{0}^{\boldsymbol{\theta}}$ are not shared with $\Gamma_{\varepsilon}^{\boldsymbol{\theta}}$.
Using the fact that any function of invariants is itself an invariant, we draw important conclusions about the structure of the system of output ODEs in Eq. (2) by analysing the linearised symmetry conditions defining parameter symmetries.
Prop. 2 (ODEs as functions of universal invariants).
Consider the system of output ODEs defined by a function $\Delta$ as in Eq. (2). Assume that this system of output ODEs has $\tilde{p}\in\{1,\ldots,p\}$ universal parameter invariants, and denote these by $I_{k},k\in\{1,\ldots,\tilde{p}\}$ which are collected in a vector $\vec{I}_{\boldsymbol{\theta}}$. Then, $\Delta$ is a function $\Phi$ of the universal invariants according to
Proof.
Let $\Gamma_{\varepsilon}^{\boldsymbol{\theta}}$ be a parameter symmetry of the system of output ODEs in Eq. (2). Then, the linearised symmetry conditions in Eq. (20) imply that the solution manifold $\Delta$ is a differential invariant of $\Gamma^{\boldsymbol{\theta}}_{\varepsilon}$, and since the same property holds for all parameter symmetries, $\Delta$ is a universal differential invariant. Accordingly, Eq. (49) follows directly from Prop. 1. ∎
Armed with Prop. 49, we understand why the standard differential algebra approach for conducting a global structural identifiability analysis (Algorithm 1) finds identifiable parameter quantities. Again, consider the case discussed previously when the system of output ODEs in Eq. (2) is defined by a function $\Delta$ which is a vector-valued multivariate polynomial. The standard differential algebra approach for elucidating structural identifiability extracts coefficients of the monomials, and by virtue of Eq. (49) these coefficients must either be a constant or a universal parameter invariant. Better still, as exemplified by the toy model previously, our notion of structural identifiability in terms of universal parameter invariants (Thm. 1 and Cor. 1) does not only yield identifiable parameter quantities, but it also allows us to characterise the parameter transformations preserving the observed outputs in the form of parameter symmetries. Next, we generalise the symmetry-based methodology for analysing structural identifiability.
@@ 3.3 The standard differential algebra approach finds universal parameter invariants
3.4 The CaLinInv-recipe: a symmetry-based approach for elucidating structural identifiability
@@
We present a recipe in three steps for elucidating structural identifiability using parameter symmetries. These three steps are captured by the acronym CaLinInv; Canonical coordinates, Linearised symmetry conditions and differential Invariants (Algorithm 2). Importantly, the CaLinInv-recipe is by no means restricted to polynomial systems of output ODEs and thus it works on arbitrary output systems. We want to emphasise that the additional information that is gained when using the CaLinInv-recipe over the differential algebra approach is the parameter symmetries or, differently put, the parameter transformations which preserve the observed outputs. In the particular case when $\Delta$ defining the system of ODEs for the outputs in Eq. (2) is composed of multivariate polynomials, the linearised symmetry conditions decompose into a system of linear equations that can be solved using Gaussian elimination.
@@
Output: The identifiable parameter quantities corresponding to universal parameter invariants and the transformations preserving these parameter quantities in the form of a family of parameter symmetries $\Gamma_{\varepsilon}^{\boldsymbol{\theta}}$.
Step 1: Canonical coordinates. The outputs $\mathbf{y}$ are canonical coordinates of full symmetries. Re-write the original system of first-order ODEs in Eq. (1) as a system of ODEs as in Eq. (2) solely depending on the observed outputs $\mathbf{y}$ and the rate parameters $\boldsymbol{\theta}$.
@@
Algorithm 2 The CaLinInv-recipe for a symmetry-based global structural identifiability-analysis.
In the case when $\Delta$ in Eq. (2) consists of rational functions of the outputs $\mathbf{y}$ and their derivatives, the linearised symmetry conditions
@@
where $M$ is a $\tilde{n}\times p$-matrix, $\boldsymbol{\chi}$ is a $p\times 1$-vector and $\mathbf{0}$ is the $\tilde{n}\times 1$-zero vector. Here, $p$ is the number of parameters, $\boldsymbol{\chi}=\left(\chi_{1},\chi_{2},\ldots,\chi_{p}\right)$ contains the parameter infinitesimals and the number of equations $\tilde{n}$ is a function of the number of monomials, e.g. “$y_{1}\dot{y}_{2}^{2}$”, in the multivariate polynomials in $\Delta$.
The first step in the CaLinInv-recipe, expressing the original system as an equivalent system solely depending on the observed outputs, is identical to the first step in the standard differential algebra approach (Algorithm 1). Thereafter, the methodologies differ where the differential algebra approach simply extracts the coefficients of the monomials and then reduces the resulting set of parameter combinations, whereas the CaLinInv-recipe solves the linearised symmetry conditions, generates parameter symmetries and calculates universal parameter invariants. In terms of outcomes, both approaches yield the identifiable parameter quantities corresponding to universal parameter invariants. However, the CaLinInv-recipe yields the family of parameter symmetries whereas the standard differential algebra approach merely yields the universal parameter invariants.
3.5 Analysing structural identifiability of glucose-insulin model with a time-dependent input
Next, we study a model of glucose-insulin regulation which was originally presented in [18]. Importantly, this model has been subject to structural identifiability analyses using the differential algebra approach [19] as well as a symmetry-based analysis focusing on full symmetries [8]. Here, we study this model by means of parameter symmetries instead and we characterise the family of parameter symmetries preserving the observed outputs.
In this model, we have two states given by $x_{1}(t)$, the glucose concentration, and $x_{2}(t)$, the insulin concentration, one known input $u(t)\geq{0}$ corresponding to the glucose entering from the digestive system, and one output $y(t)$ corresponding to a glucose measurement. In total, there are five parameters
@@ 3.5 Analysing structural identifiability of glucose-insulin model with a time-dependent input
where the first four encode first-order reaction rates while the last parameter corresponds to the volume of blood extracted during glucose measurement and the original system of ODEs is given by
$\begin{split}\dot{x_{1}}&=u+p_{1}x_{1}-p_{2}x_{2},\\ \dot{x_{2}}&=p_{3}x_{2}+p_{4}x_{1},\\ y&=\frac{x_{1}}{V_{p}}.\end{split}$ (53)
@@
$V_{p}\ddot{y}-V_{p}(p_{1}+p_{3})\dot{y}+V_{p}(p_{1}p_{3}+p_{2}p_{4})y+p_{3}u-\dot{u}=0,$ (54)
and the linearised symmetry condition is given by
$\begin{split}[-V_{p}\dot{y}+V_{p}p_{3}y]\chi_{p_{1}}+[V_{p}p_{4}y]\chi_{p_{2}}+[-V_{p}\dot{y}+V_{p}p_{1}y]\chi_{p_{3}}+u\chi_{p_{3}}+[V_{p}p_{2}y]\chi_{p_{4}}&\\ [\ddot{y}-(p_{1}+p_{3})\dot{y}+(p_{1}p_{3}+p_{2}p_{4})y]\chi_{V_{p}}&=0.\\ \end{split}$ (55)
Therefore, the linearly independent set of coefficients is those relating to $\{\ddot{y},\dot{y},y,u\}$. The coefficient of $\ddot{y}$ yields that $\chi_{V_{p}}=0$ and hence $V_{p}$ is structurally identifiable. Moreover, the coefficient of the input $u$ yields that $\chi_{p_{3}}=0$ and hence $p_{3}$ is also structurally identifiable. By substituting in $\chi_{V_{p}}=\chi_{p_{3}}=0$ into Eq. (55), we obtain
$\dot{y}\left(-V_{p}\chi_{p_{1}}\right)+y\left(V_{p}p_{3}\chi_{p_{1}}+V_{p}p_{4}\chi_{p_{2}}+V_{p}p_{2}\chi_{p_{4}}\right)=0.$ (56)
The coefficient of $\dot{y}$ yields that $\chi_{p_{1}}=0$ and hence $p_{1}$ is structurally identifiable. Lastly, the coefficient of $y$ together with the conclusion that $\chi_{p_{1}}=0$ yields
@@
In conclusion, the parameters $p_{1}$, $p_{3}$ and $V_{p}$ are globally structurally identifiable. The parameters $p_{2}$ and $p_{4}$ are globally structurally unidentifiable whereas their product $p_{2}p_{4}$ is globally structurally identifiable.
The glucose-insulin model considered here consists of two first-order ODEs, one output equation and five rate parameters, and for such a small model we can calculate the universal parameter invariants and parameter symmetries using simple calculations by hand. For larger models with more equations and parameters, the linearised symmetry conditions decompose into a matrix system which can be solved using Gaussian elimination. We next demonstrate this fact by analysing the global structural identifiability of a more complicated model.
3.6 Analysing structural identifiability of an epidemiological model
$\begin{split}\dfrac{\mathrm{d}S}{\mathrm{d}t}&=c-\beta SI-\mu_{S}S,\\ \dfrac{\mathrm{d}E}{\mathrm{d}t}&=(1-\upsilon)\beta SI-\delta E-\mu_{E}E,\\ \dfrac{\mathrm{d}I}{\mathrm{d}t}&=\upsilon\beta SI+\delta E-\mu_{I}I.\\ \end{split}$ (63)

### v3

@@ 4.1 Analysing the structural identifiability of a toy model using parameter symmetries
This toy example illustrates that the standard differential algebra approach and the approach based on parameter symmetries arrive at the same conclusions regarding the identifiable parameter quantities in the case where the locally and globally structurally identifiable parameter quantities are the same. Additionally, the symmetry-based approach also yields the family of parameter symmetries corresponding to the parameter transformations that preserve the observed outputs.
Next, we analyse the structural identifiability of a linear model, for which the locally and globally identifiable parameters are different, by means of the standard differential algebra approach and the approach based on parameter symmetries.
4.2 Analysing structural identifiability of a linear model where the locally and globally identifiable parameters are different
Let $u(t)$ and $v(t)$ denote the concentrations of two chemical species depending on time $t$ which satisfy the coupled system of two linear ODEs
[…]
where $a,b,c$ are three positive rate constants. Furthermore, assume that we observe the first species, i.e. $y(t)=u(t)$, yielding the following model for the output
$0=\dfrac{\mathrm{d}^{2}y}{\mathrm{d}t^{2}}-(a+c)\dfrac{\mathrm{d}y}{\mathrm{d}t}+(ac)y\,.$ (40)
We first apply the standard differential algebra approach for analysing the structural identifiability of this model. First, we see that the parameter $b$ does not appear in the output equation in Eq. (40), and hence $b$ is structurally non-identifiable. Next, we extract the coefficients in front of $\{\mathrm{d}^{2}y/\mathrm{d}t^{2},\mathrm{d}y/\mathrm{d}t,y\}$ resulting in the set $\{1,\mathcal{A},\mathcal{B}\}$ where $\mathcal{A}=a+c$ is the sum and $\mathcal{B}=ac$ is the product of the parameters $a$ and $c$. Clearly, the globally structurally identifiable parameter quantities correspond to the sum $\mathcal{A}$ and the product $\mathcal{B}$, whereas the individual parameters $a$ and $c$ are not globally structurally identifiable. This can be seen by equating $a=\mathcal{A}-c$ with $a=\mathcal{B}/c$ which yields a quadratic equation for the parameter $c$:
$c^{2}-c\mathcal{A}+\mathcal{B}=0\,.$ (41)
The two solutions of this equation are given by
$c=\frac{1}{2}\left(\mathcal{A}\pm\sqrt{\mathcal{A}^{2}-4\mathcal{B}}\right)\,$, (42)
and thus the two sets of parameter values that yield the same outputs are given by
$(a,c)=\left(\dfrac{2\mathcal{B}}{\left(\mathcal{A}\pm\sqrt{\mathcal{A}^{2}-4\mathcal{B}}\right)},\frac{1}{2}\left(\mathcal{A}\pm\sqrt{\mathcal{A}^{2}-4\mathcal{B}}\right)\right)\,.$ (43)
Thus, the parameters $a$ and $c$ are only locally structurally identifiable. Furthermore, since parameter symmetries act on parameters locally in an infinitesimal fashion, it is not surprising that their universal parameter invariants indeed correspond to the locally identifiable parameter quantities. This is clearly illustrated using the symmetry-based method.
We now analyse the structural identifiability of the output system in Eq. (40) using parameter symmetries. The linearised symmetry condition is given by
$0=-(\chi_{a}+\chi_{c})\dfrac{\mathrm{d}y}{\mathrm{d}t}+(c\chi_{a}+a\chi_{c})y\,.$ (44)
In order for a parameter to be locally structurally identifiable, we require its infinitesimal to be identically equal to zero. Since the infinitesimal $\chi_{b}$ does not appear in the above equation, it follows that $b$ is non-identifiable. Decomposing Eq. (44) with respect to the linearly independent set $\{\mathrm{d}y/\mathrm{d}t,y\}$ yields
$\chi_{a}=-\chi_{c}\,,\quad\chi_{a}=-\left(\dfrac{a}{c}\right)\chi_{c}\,.$ (45)
Equating these two equations yields
$\left(1-\dfrac{a}{c}\right)\chi_{c}=0\,$,
and when $a\neq{c}$ it follows that $\chi_{c}=0$ and thus $c$ is locally structurally identifiable. Since $\chi_{a}=-\chi_{c}$, we have that $\chi_{a}=0$ and thus $a$ is also locally structurally identifiable. Thus, when $a\neq{c}$ both $a$ and $c$ are locally structurally identifiable.
When $a=c$ we see that the output ODE in Eq. (40) becomes $0=\mathrm{d}^{2}y/\mathrm{d}t^{2}-2a(\mathrm{d}y/\mathrm{d}t)+a^{2}y$ and the corresponding linearised symmetry condition becomes $0=-2\chi_{a}(\mathrm{d}y/\mathrm{d}t)+2a\chi_{a}y$. By separating the last linearised symmetry condition with respect to $\{\mathrm{d}y/\mathrm{d}t,y\}$, it decomposes into the single constraint $\chi_{a}=0$ implying that $a$ is locally structurally identifiable. Since $a=c$, $c$ must also be locally structurally identifiable.
In conclusion, this example highlights the key difference between the standard differential algebra approach to structural identifiability and the parameter-symmetry-based approach. The former identifies globally identifiable parameter combinations, whereas the latter reveals locally structurally identifiable parameter quantities. Next, we generalise the parameter-symmetry-based approach for analysing local structural identifiability.
@@
4.3 A parameter-symmetry-based approach for elucidating local structural identifiability
@@
We present a recipe in three steps for elucidating local structural identifiability using parameter symmetries. These three steps are captured by the acronym CaLinInv; Canonical coordinates, Linearised symmetry conditions and differential Invariants (Alg. 1). Importantly, the CaLinInv recipe is by no means restricted to polynomial systems of output ODEs and thus it works on arbitrary output systems. We want to emphasise that the additional information that is gained when using the CaLinInv recipe over the differential algebra approach is the parameter symmetries or, differently put, the parameter transformations which preserve the observed outputs. In the particular case when $\Delta$ defining the system of ODEs for the outputs in Eq. (4) is composed of multivariate polynomials, the linearised symmetry conditions decompose into a system of linear equations that can be solved using Gaussian elimination.
@@
Output: The locally identifiable parameter quantities corresponding to universal parameter invariants and the transformations preserving these parameter quantities in the form of a family of parameter symmetries $\Gamma_{\varepsilon}^{\boldsymbol{\theta}}$.
Step 1: Canonical coordinates. The outputs $\mathbf{y}$ are canonical coordinates of parameter symmetries. Re-write the original system of first-order ODEs in Eq. (3) as a system of ODEs as in Eq. (4) solely depending on the observed outputs $\mathbf{y}$ and the parameters $\boldsymbol{\theta}$.
@@
Algorithm 1 The CaLinInv-method for a symmetry-based local structural identifiability-analysis.
In the case where $\Delta$ in Eq. (4) consists of rational functions of the outputs $\mathbf{y}$ and their derivatives, the linearised symmetry conditions
@@
where $M$ is a $\tilde{n}\times p$-matrix, $\boldsymbol{\chi}$ is a $p\times 1$-vector and $\mathbf{0}$ is the $\tilde{n}\times 1$-zero vector. Here, $p$ is the number of parameters and $\boldsymbol{\chi}=\left(\chi_{1},\chi_{2},\ldots,\chi_{p}\right)$ contains the parameter infinitesimals.
The first step in the CaLinInv recipe, expressing the original system as an equivalent system solely depending on the observed outputs, is identical to the first step in the standard differential algebra approach. Thereafter, the methodologies differ where the differential algebra approach simply extracts the coefficients of the monomials and then reduces the resulting set of parameter combinations, whereas the CaLinInv recipe solves the linearised symmetry conditions, generates parameter symmetries and calculates universal parameter invariants. In terms of outcomes, both approaches yield the identifiable parameter quantities corresponding to universal parameter invariants. However, the CaLinInv recipe yields the family of parameter symmetries whereas the standard differential algebra approach merely yields the universal parameter invariants.
4.4 Analysing local structural identifiability of a glucose-insulin model with a time-dependent input
Next, we study a model of glucose-insulin regulation that was originally presented in [20]. Importantly, this model has been subject to structural identifiability analyses using the differential algebra approach [21] as well as a symmetry-based analysis focusing on full symmetries [8]. Here, we study this model by means of parameter symmetries instead and we characterise the family of parameter symmetries preserving the observed outputs.
Step 1: Canonical coordinates
In this model, we have two states given by $x_{1}(t)$, glucose concentration, and $x_{2}(t)$, insulin concentration, one known input $u(t)\geq{0}$ corresponding to glucose entering the digestive system, and one output $y(t)$ corresponding to a glucose measurement. In total, there are five parameters
@@ 4.4 Analysing local structural identifiability of a glucose-insulin model with a time-dependent input
where the first four encode first-order reaction rates while the last parameter corresponds to the volume of blood extracted during glucose measurement, and the original system of ODEs is given by
$\begin{split}\dfrac{\mathrm{d}x_{1}}{\mathrm{d}t}&=u+p_{1}x_{1}-p_{2}x_{2},\\ \dfrac{\mathrm{d}x_{2}}{\mathrm{d}t}&=p_{3}x_{2}+p_{4}x_{1},\\ y&=\frac{x_{1}}{V_{p}}.\end{split}$ (49)
@@
$V_{p}\dfrac{\mathrm{d}^{2}y}{\mathrm{d}t^{2}}-V_{p}(p_{1}+p_{3})\dfrac{\mathrm{d}y}{\mathrm{d}t}+V_{p}(p_{1}p_{3}+p_{2}p_{4})y+p_{3}u-\dfrac{\mathrm{d}u}{\mathrm{d}t}=0\,.$ (50)
Step 2: Linearised symmetry condition
The linearised symmetry condition is given by
[…]
Therefore, the linearly independent set of coefficients is those relating to $\{\mathrm{d}^{2}y/\mathrm{d}t^{2},\mathrm{d}y/\mathrm{d}t,y,u\}$. The coefficient of $\mathrm{d}^{2}y/\mathrm{d}t^{2}$ shows that $\chi_{V_{p}}=0$ and therefore $V_{p}$ is locally structurally identifiable. Moreover, the coefficient of the input $u$ yields that $\chi_{p_{3}}=0$ and therefore $p_{3}$ is also structurally identifiable. By substituting $\chi_{V_{p}}=\chi_{p_{3}}=0$ into Eq. (51), we obtain
$\dfrac{\mathrm{d}y}{\mathrm{d}t}\left(-V_{p}\chi_{p_{1}}\right)+y\left(V_{p}p_{3}\chi_{p_{1}}+V_{p}p_{4}\chi_{p_{2}}+V_{p}p_{2}\chi_{p_{4}}\right)=0.$ (52)
The coefficient of $\mathrm{d}y/\mathrm{d}t$ yields that $\chi_{p_{1}}=0$ and hence $p_{1}$ is locally structurally identifiable. Lastly, the coefficient of $y$ together with the conclusion that $\chi_{p_{1}}=0$ yields
@@
In conclusion, the parameters $p_{1}$, $p_{3}$ and $V_{p}$ are locally structurally identifiable. The parameters $p_{2}$ and $p_{4}$ are structurally non-identifiable whereas their product $p_{2}p_{4}$ is locally structurally identifiable.
The glucose-insulin model considered here consists of two first-order ODEs, one output equation and five parameters, and for such a small model we can calculate the universal parameter invariants and parameter symmetries using simple calculations by hand. For larger models with more equations and parameters, the linearised symmetry conditions decompose into a matrix system which can be solved using Gaussian elimination. We next demonstrate this fact by analysing the local structural identifiability of a more complicated model.
4.5 Analysing local structural identifiability of an epidemiological model
We study an SEI model of the epidemiology of tuberculosis [16]. This model consists of three states: $S(t)$ corresponds to the susceptible population; $E(t)$ corresponds to the exposed population; and $I(t)$ corresponds to the infected population. Moreover, this model has seven parameters: $c$ is the birth rate; $\beta$ is the transmission rate; $\upsilon$ is the probability of primary infection; $\delta$ is the reactivation rate and $\mu_{S}$, $\mu_{E}$ and $\mu_{I}$ are death rates. All parameters are assumed to be positive.
We implement the CaLinInv recipe (Alg. 1) to analyse structural identifiability. Importantly, we compare the outcomes of these calculations to those obtained through the standard differential algebra approach. Also, we generate and visualise parameter symmetries of this model. The underlying semi-manual calculations were conducted using the open-source symbolic solver SymPy [22], and they are presented in the appendices. Details and relevant scripts are available at the public GitHub-repository associated with this project: https://github.com/JohannesBorgqvist/symmetries_and_structural_identifiability.
Step 1: Canonical coordinates
The original system of ODEs is given by
[…]
We consider two outputs given by a proportion of the exposed population and a proportion of the infected population, where these proportions are encoded by the parameters $k_{E}$ and $k_{I}$, respectively. Then, the two observed outputs denoted by $y_{E}$ and $y_{I}$ are given by